# From Perception to Reasoning: Deep Thinking Empowers Multimodal Large Language Models

### v1

@@ 6.3 Omnimodal Reasoning
The GPT-4o (mini) model demonstrates significant potential of large models in perceiving and processing multimodal data; however, it lacks chain-of-though reasoning capabilities. Currently, multimodal CoT reasoning primarily focuses on bimodal combinations such as text-image (liu2025visual; li2024streetviewllm), text-video(fei2024video; himakunthala2023let; zhang2024worldqa), or text-audio(du2024cot; xie2025audio), with relatively limited modality coverage, making it insufficient to meet more complex and general reasoning scenarios.
@@
In this domain, R1-Omini(zhao2025r1) achieved full-modality expansion of CoT reasoning in emotion recognition tasks and was the first to introduce verifiable reward-based reinforcement learning into full-modality large language models for emotion recognition, where text, visual, and audio modalities all played crucial roles.
@@ 6.4 Efficient Reasoning
Regarding reasoning efficiency evaluation, MME-CoT(jiang2025mme) introduces two metrics—relevance and reflection quality—to measure the effectiveness of multimodal chain-of-thought reasoning. To address inefficiencies caused by verbose texts in multimodal reasoning, Shen et al.(shen2025efficient) propose the Heima framework, which aims to enable efficient hidden reasoning. This approach fine-tunes the Heima encoder by encoding each reasoning chain as a single token and incorporates interpretable prompt tokens in the decoder to decode the hidden representations encapsulated in the reasoning tokens. Lu et al.(lu2025prolonged) propose a deterministic adaptive reasoning framework called CAR, which dynamically switches between System 1 and System 2 reasoning based on the model’s perplexity. CAR initially employs System 1 reasoning to generate concise answers and transitions to System 2 for deeper reasoning when the model exhibits high perplexity toward the initial answer.
Wang et al.(wang2025harnessing) and Feng et al.(feng2025efficient) conduct systematic surveys on the reasoning efficiency of large language models and propose a series of efficient reasoning strategies applicable to both post-training and inference stages, some of which have been successfully applied in multimodal large models.

### v2

@@ 6.3 Omnimodal Reasoning
The GPT-4o (mini) model demonstrates significant potential of large models in perceiving and processing multimodal data; however, it lacks chain-of-though reasoning capabilities. Currently, multimodal CoT reasoning primarily focuses on bimodal combinations such as text-image (Liu et al., 2025e; Li et al., 2024c), text-video(Fei et al., 2024; Himakunthala et al., 2023; Zhang et al., 2024g), or text-audio(Du et al., 2024; Xie et al., 2025), with relatively limited modality coverage, making it insufficient to meet more complex and general reasoning scenarios.
@@
In this domain, R1-Omini(Zhao et al., 2025b) achieved full-modality expansion of CoT reasoning in emotion recognition tasks and was the first to introduce verifiable reward-based reinforcement learning into full-modality large language models for emotion recognition, where text, visual, and audio modalities all played crucial roles.
@@ 6.4 Efficient Reasoning
Regarding reasoning efficiency evaluation, MME-CoT(Jiang et al., 2025a) introduces two metrics—relevance and reflection quality—to measure the effectiveness of multimodal chain-of-thought reasoning. To address inefficiencies caused by verbose texts in multimodal reasoning, Shen et al.(Shen et al., 2025a) propose the Heima framework, which aims to enable efficient hidden reasoning. This approach fine-tunes the Heima encoder by encoding each reasoning chain as a single token and incorporates interpretable prompt tokens in the decoder to decode the hidden representations encapsulated in the reasoning tokens. Lu et al.(Lu et al., 2025) propose a deterministic adaptive reasoning framework called CAR, which dynamically switches between System 1 and System 2 reasoning based on the model’s perplexity. CAR initially employs System 1 reasoning to generate concise answers and transitions to System 2 for deeper reasoning when the model exhibits high perplexity toward the initial answer.
Wang et al.(Wang et al., 2025c) and Feng et al.(Feng et al., 2025) conduct systematic surveys on the reasoning efficiency of large language models and propose a series of efficient reasoning strategies applicable to both post-training and inference stages, some of which have been successfully applied in multimodal large models.